Shift report 20000820 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Aug 20 2000 - 08:00:27 EDT

  • Next message: Konstantin Olchanski: "pii4 status"

    ------------------------------------------------------------
    Date: Sun Aug 20 08:00:26 2000, Shift: 00:00-08:00
    Supervisor: sjs, with: sjs, ej
    Log book#4 pages 128 - 129
    
    Shift Summary:
    ------------------------------------------------------------
    1.) No beam until 4:30, then (finally) a good store.
    2.) Runs 2355-2357,  many collisions (too many sequences
    to complete BrNtuple run during shift, even for the first
    1 hr run (2355)).
    Moved Si/Tile monitoring program to operator account. 
    Instructions for use placed in "BRAHMS Experiment Information"
    book next to BrNtuple program instructions.  This should
    be run for each significant data run.
    EJ discovered temperature in FE Hut at 77 deg.--
    The new termostat has both weekday and weekend setting--
    the weekend defaults apparently were never changed.  Reset
    to maintain constant 70 deg.   (Some evidence that the
    higher temperature might have caused significant problems
    for some of the Si detectors.) 
    3.)  No new detector problems discovered.  However, we
    were unable to print out the "Shifters Checklist".  Has
    someone moved this???
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Aug 20 2000 - 08:07:20 EDT